Circular No. 5951

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1994D IN NGC 4526 R. J. Cumming, Royal Greenwich Observatory (RGO); W. P. S. Meikle,

Imperial College; and T. R. Geballe, Joint Astronomy Center (JAC), report: “Near-infrared spectra (resolution 800–1200 km/s) of SN 1994D were obtained by J. V. Wall (RGO), C. R. Jenkins (RGO), Geballe, and D. M. Walther (JAC) with the U.K. Infrared Telescope (+ CGS4) on Mar. 12.52 (J, H, K bands) and 13.51 UT (K only). The J-band spectrum shows a wide absorption feature centered at 1.052 μm and extending blueward as far as 1.036 μm. If this is due to a P-Cyg profile of He I 1.083-μm, it implies ejecta velocities as high as 14 000 km/s. On Mar. 12.52, the Kband spectrum showed absorption features centred at 2.30, 2.33, 2.36, and 2.39 μm, all having a depth of 8–9 percent of the continuum. We identify these features with first-overtone CO absorption at the redshift of NGC 4526. The 2–0 band head at 2.300 μm is unresolved at a resolution of 850 km/s. The spectrum of Mar. 13.51 shows the 2–0 band at approximately the same strength and the band head still unresolved, but the other CO bands have weakened considerably. An optical spectrum taken using ISIS on the William Herschel Telescope on Mar. 10.14 by L. J. Smith (University College London), M. Pettini and D. L. King (RGO), and C. Martín (Isaac Newton Group) shows no evidence for narrow emission or absorption in Hα associated with the supernova. We judge that the origin of the CO features must be intrinsic to the supernova system. This is supported by the change in their visibility. The early appearance and narrowness of the features imply that the CO lay above the ejecta.”

T. Shanks and S. M. Croom, University of Durham; and N. R. Tanvir, Insititute of Astronomy, report: “We obtained provisional photometry (±0.01) of SN 1994D using the Jacobus Kapteyn 1.0-m telescope: Mar. 10.16 UT, U = 13.4, B = 13.7, V = 13.7, R = 13.6, I = 13.7; 11.17, 12.8, 13.3, 13.3, 13.2, 13.2. The B − V color is close to the mean for type-Ia supernovae at maximum. The supernova appears 3′′–4′′ to the north of a dust lane in this early-type galaxy. These observations are consistent with a Virgo type-Ia supernova nearing maximum in the next few days.”

P. M. Kilmartin and A. C. Gilmore, Mount John, report the following equinox 2000.0 position for SN 1994D: α = 12^{h}34^{m}02^{s}.37, δ = +7^{o}42′04′′.7 (corresponding 1950.0 position is α = 12^{h}31^{m}29^{s}.99, δ = +7^{o}58′36′′.5).

1994 March 14 Daniel W. E. Green